Ok, I have two questions here.....and Im really stuck.. I would be very grateful if you found the solutions.
1. A parallelogram WXYZ has three vertices at W(2,1),X(-1,5),Y(-3,3).
a) Find the Midpoint of WY
b) Use this information to find the Co-ordinates of Z

Answer by Alan3354(69443)

1. A parallelogram WXYZ has three vertices at W(2,1),X(-1,5),Y(-3,3).
a) Find the Midpoint of WY
Find the x and y separately. Use the average of x and y
For x: (2 - 3)/2 = -1/2
For y: (1 + 3)/2 = 2
--> M(-1/2,2)
------------
b) Use this information to find the Co-ordinates of Z
The midpoint M is also the midpoint of the other diagonal XZ
X is (-1,5) and M is (-1/2,2)
The x value increases 1/2 to the midpoint, so it increases 1/2 again to point X.
-1/2 + 1/2 = 0
----------
The y value decreases 3 to the midpoint, so it decreases 3 again to point X.
2 - 3 = -1
--> X (0,-1)